# I-INR: Iterative Implicit Neural Representations

### v2

@@ Abstract
Implicit Neural Representations (INRs) have revolutionized signal processing and computer vision by modeling signals as continuous, differentiable functions parameterized by neural networks. However, their inherent formulation as a regression problem makes them prone to regression to the mean, limiting their ability to capture fine details, retain high-frequency information, and handle noise effectively. To address these challenges, we propose Iterative Implicit Neural Representations (I-INRs) a novel plug-and-play framework that enhances signal reconstruction through an iterative refinement process. I-INRs effectively recover high-frequency details, improve robustness to noise, and achieve superior reconstruction quality. Our framework seamlessly integrates with existing INR architectures, delivering substantial performance gains across various tasks. Extensive experiments show that I-INRs outperform baseline methods, including WIRE, SIREN, and Gauss, in diverse computer vision applications such as image restoration, image denoising, and object occupancy prediction.
Figure 1: Effectiveness of the proposed methods across multiple tasks, including image fitting, Super Resolution (2$\times$ scale), and denoising, compared to the baseline representative INR method (SIREN [39]). Our novel Iterative-INR method (I-SIREN) consistently improves detail preservation, fidelity, and high-frequency reconstruction across all tasks, outperforming the baseline.
@@ 1 Introduction
Implicit Neural Representations (INRs) have emerged as a transformative approach in signal representation, shifting from traditional discrete grid-based methods to continuous coordinate-based models [39]. By leveraging neural networks, typically multi-layer perceptrons (MLPs), INRs map spatial or temporal coordinates directly to signal attributes such as pixel intensity, colour, or 3D occupancy. This continuous representation inherently offers resolution independence, compact encoding, and seamless interpolation, making INRs highly versatile across a wide range of domains including computer vision [5, 35, 39, 24, 18], and beyond [42, 37, 49, 31]. Initially introduced for tasks like 3D shape reconstruction [24, 26, 41] and novel view synthesis [17], INRs have since been applied to diverse applications including image fitting [23], super-resolution [9, 34], and solving inverse problems [3, 44, 5].
Despite advancements, INRs continue to face substantial challenges in capturing high-frequency details, maintaining robustness against noise, and handling incomplete data [36]. Typically optimized with L1 or L2 loss [39], INRs exhibit a bias toward low-frequency components due to the regression to the mean effect, leading to the loss of high-frequency details and reduced noise robustness [9]. Recent works have tried to solve this problem through positional encoding techniques  [43] or by incorporating new activation functions  [36, 39, 23]. Positional encoding techniques [43] incorporate various orthogonal Fourier [11], while periodic activations, such as sine functions, allow MLPs to better model high-frequency signals compared to conventional ReLU activations [39]. However, these methods are often based on the assumption of clean, noise-free data, which limits their applicability in real-world scenarios where noise and missing information are prevalent [36]. Advanced activation functions like complex Gabor wavelets have been proposed to enhance noise robustness [36], but these approaches still fall short when handling the complexities of challenging, real-world datasets. Despite these efforts, there remains significant potential for further improving high-frequency detail preservation, representational accuracy, and reconstruction quality in INRs.
@@
To address these limitations, drawing inspiration from iterative models [9, 33, 15], we introduce Iterative Implicit Neural Representations (I-INRs), a plug-and-play framework that integrates seamlessly with existing implicit architectures. Unlike traditional single-shot INRs, I-INRs reconstruct signals progressively over multiple iterations, refining details, enhancing reconstruction quality, and improving noise robustness as shown in Figure 1. Our experiments demonstrate that I-INRs deliver superior reconstructions with enhanced detail, generalization, and minimal computational overhead, outperforming single-shot methods like WIRE [36], SIREN [39], and Gauss [30] across tasks such as image fitting, Super-Resolution (SR), denoising, and 3D occupancy reconstruction.
Our main contributions are.
@@ 2 Related Work
INRs [39] map coordinates to signal values using neural networks, eliminating the need for discretization and thus offering flexibility in applications such as image, video processing [14, 5, 46, 4, 7], and 3D rendering [24, 32, 13]. However, INRs often suffer from spectral bias [29, 43], struggling to capture high-fidelity details of complex signals. This limitation affects applications requiring fine-grained detail, such as image SR, 3D reconstruction, and scene modeling. To address this challenge, various approaches have been proposed, including spatial encoding with frequency-based representations, polynomial decomposition [28, 38], and high-pass filtering [11], all of which emphasize high-frequency components. Beyond encoding techniques, the choice of activation functions also plays a crucial role in mitigating spectral bias. Standard activations like ReLU fail to capture high-frequency components due to their limited representational capacity [39]. To overcome this, alternative activation functions, such as sinusoidal [39] and Gaussian [30] activations, have been proposed. Additionally, approaches like WIRE [36] and FINER [23] introduce advanced activation strategies to enhance spectral bias tuning, with WIRE [36] leveraging Gabor wavelets and FINER [23] utilizing variable-periodic functions. INCODE [19] uses a task-specific pre-trained model that dynamically adjusts key parameters of activation function to enhance reconstruction performance. DINER [47] utilizes a hash map to unevenly map input coordinates onto a feature vector, optimizing the spatial frequency distribution and enabling faster, more precise reconstructions. While advanced encoding strategies and activation functions help mitigate spectral bias, they may not fully address the challenges of wide-bandwidth signals and large-scale data  [21, 35]. To scale INRs for wide-bandwidth signals, hierarchical and multi-scale designs like Progressive Implicit Networks (PINs) [21] and MINER [35] leverage hierarchical encoding and Laplacian pyramids, enabling efficient multi-scale detail capture while reducing computational overhead.
@@
In contrast with single shot models, iterative models, such as diffusion models [15, 33, 9], generate high-quality samples by inverting known degradation processes. These models are widely used for diverse image processing tasks, including image restoration and translation [34, 33, 9, 16, 8, 6]. Furthermore, diffusion models have also been applied to 3D reconstruction tasks from text prompts or 2D images [27, 22]. Inspired by the strong detail reconstruction capabilities of iterative models, we propose a novel INR framework that reconstructs signals implicitly over multiple steps. Our method is a flexible, plug-and-play approach that can be integrated into any existing INR framework. However, for the scope of this paper, we specifically focus on WIRE [36], Gauss [30], and SIREN [39].
@@
Figure 2: (a) The proposed architecture of the I-INR model. The framework consists of a Base INR, which can be any baseline architecture, denoted as Backbone, such as SIREN [39], WIRE [36], or Gauss [30]. Additionally, an Auxiliary Network incorporates feedback to refine representations, while a projection layer integrates features. The final output is obtained by combining the outputs of the Base INR and the projection layer, enhancing expressivity and reconstruction quality. (b) Iterative reconstruction processes of the proposed I-INR framework.
@@ 3 Proposed Method
In this section, we first provide an overview of INRs, followed by a detailed explanation of our proposed method, I-INR.
3.1 Preliminary: Standard INR
Let $g:X\subseteq\mathbb{R}^{p}\rightarrow Y\subseteq\mathbb{R}^{q}$ be a target function mapping $p$-dimensional inputs (such as pixel positions in a 2D image or volumetric domains in 3D scenes) to $q$-dimensional outputs (such as color or density values), so that $y=g(x)$ for ${}{x\in X}$. An INR [39] is a learnable neural function $f_{\theta}:X\rightarrow Y$, parameterized by $\theta$, aiming to approximate $\mathcal{I}(x)$ across $X$ [36]. The objective is to find $\theta$ such that:
$f_{\theta}(x)\approx\mathcal{I}(x),$ (1)
where $\mathcal{I}(x)$ represents an image function with pixel coordinates as $x$. A typical INR is trained by reducing the signal error in some $L_{p}$ metric,
$\min_{\theta}\mathbb{E}_{x,\mathcal{I}(x)}\|f_{\theta}(x)-\mathcal{I}(x)\|_{p}\approx\min_{\theta}\sum_{i}\|f_{\theta}(x^{i})-\mathcal{I}(x^{i})\|_{p},$ (2)
where $p=2$, the optimal minimum mean squared error (MMSE) solution is the conditional expectation as in Eq. (3)
@@ 3.1 Preliminary: Standard INR
This expectation leads to a signal that represents the average of all plausible reconstructions, often resulting in the loss of fine details. Such an approach limits the ability to capture high-frequency components and makes the reconstruction susceptible to noise. To overcome these limitations, we propose an iterative reconstruction approach that refines the signal progressively over multiple steps as formulated in Section 3.2. This iterative method enables a more precise reconstruction, enhancing frequency details and improving robustness to noise.
3.2 Iterative Implicit Neural Representations: I-INRs
I-INR leverages InDI [9] for its iterative setup. It reconstructs the signal in multiple steps, improving the reconstruction quality at each step. The function $g(x)_{t}$ at any time step $t$ can be defined as:
$g(x)_{t}=\mathcal{I}(x)(1-t)+\mathcal{Z}\,t,\quad t\in[0,1],$ (4)
where $\mathcal{Z}$ is sampled from a known distribution with the same dimensions as $\mathcal{I}(x)$. The process in Eq. (4) linearly interpolates between $\mathcal{I}(x)$ and $\mathcal{Z}$ using the time step $t$. At $t=0$, $g(x)_{0}$ equates to $\mathcal{I}(x)$, while at $t=1$, $g(x)_{1}$ is $\mathcal{Z}$. Thus, $g(x)_{t}$ smoothly transitions from $\mathcal{I}(x)$ to $\mathcal{Z}$ as $t$ moves from 0 to 1, providing a continuous interpolation between the two known states as shown in Figure 2(b).
During the iterative restoration process, we start from $\mathcal{Z}$ at $t=1$ and move toward the final reconstruction gradually. Let, $g(x)_{s}$ and $g(x)_{t}$ represent some intermediate states such that $s\leq t$. Then the conditional mean is given by:
$\mathbb{E}(g(x)_{s}\mid g(x)_{t})=\left(1-\frac{s}{t}\right)\mathbb{E}(g(x)_{0}\mid g(x)_{t})+\frac{s}{t}g(x)_{t}.$ (5)
When, $s=t-\delta$, then the estimated state $\hat{g}(x)_{t-\delta}$ is given by the Eq. 6 as:
$\begin{split}\hat{g}(x)_{t-\delta}&=\mathbb{E}(g(x)_{t-\delta}\mid g(x)_{t})\\ &=\frac{\delta}{t}\mathbb{E}(g(x)_{0}\mid\hat{g}(x)_{t})+\left(1-\frac{\delta}{t}\right)\hat{g}(x)_{t}.\end{split}$ (6)
Following Eq. (6), the reconstruction process begins at $g(x)_{1}=\mathcal{Z}$, with $t=1$. The final reconstruction is approached iteratively, taking steps of size $\delta$, by estimating the signal at each intermediate step.
Furthermore, taking the limit $\delta\to 0$ in Eq. (6) leads to the following ordinary differential equation (ODE)  [40, 9]:
$\frac{dg(x)_{t}}{dt}$ $=\lim_{\delta\to 0}\frac{g(x)_{t}-g(x)_{t-\delta}}{\delta}$ (7)
$=\frac{g(x)_{t}-\mathbb{E}(g(x)_{0}\mid g(x)_{t})}{t}.$
Here, the term $g(x)_{t}-\mathbb{E}[g(x)_{0}\mid g(x)_{t}]$ represents the residual error correction in the reconstruction process, where $\frac{1}{t}$ scales the update, allowing larger corrections in the early stages when $t$ is large, while finer adjustments take place as $t\to 0$, ensuring a smooth and gradual refinement toward $g(x)_{0}$ effectively preventing over-smoothing.
I-INR Training: Our proposed method aims to generate a progressively refined version of the image at each time step. To achieve this, we estimate a reconstructed version of the signal at each intermediate step $t$, specifically $\mathbb{E}(g(x)_{0}\mid g(x)_{t})$. This is achieved by training our neural function to predict the signal itself at any intermediate time step.
Given ${x},{I(x)}\sim p({x},{I(x)})$, $\mathcal{Z}\sim p(\mathcal{Z})$, and $t\sim p(t)$, the ideal objective function for this process is defined as:
$\min_{\theta}\mathbb{E}_{({x},{I(x)}),{\mathcal{Z}}}\mathbb{E}_{t}\|f_{\theta}({g(x)}_{t},{x},t)-{I(x)}\|_{2},$ (8)
where,
${g}(x)_{t}=\mathcal{I}(x)(1-t)+\mathcal{Z}t.$
To follow the regularity requirement mentioned in [9] for the iterative procedure, we add small noise to the intermediate latent state. The model with noise perturbation is given by:
$\tilde{g}_{t}(x)=I(x)(1-t)+\mathcal{Z}t+\epsilon\cdot n\cdot t,$ (9)
where, $\epsilon$ denotes a small constant that ensures a controlled amount of noise, and $n$ is a random variable sampled from a standard normal distribution $\mathcal{N}(0,1)$. Hence, our training objective becomes,
$\min_{\theta}\mathbb{E}_{({x},{I(x)}),{\mathcal{Z}}}\mathbb{E}_{t}\mathbb{E}_{n}\|f_{\theta}(\tilde{g}(x)_{t},{x},t)-{I(x)}\|_{2}.$ (10)
The detailed training procedure is provided in Algo. 1.
I-INR Reconstruction: During reconstruction, we start from an initial state $\mathcal{Z}$ and iteratively move towards the final reconstruction over $steps=\frac{1}{\delta}$. Typically, the known function $\mathcal{Z}$ is a simple initial state. Thus, we begin from $\mathcal{Z}$ and iteratively enhance it over time. The iterative implicit reconstruction process is outlined in Algo. 2.
Algorithm 1 I-INRs Model Training $f_{\theta}$
$\epsilon$
$(x,I(x))\sim p(x,I(x))$
$\mathcal{Z}\sim p(\mathcal{Z})$
repeat
     $t\sim\mathcal{U}(0,1)$
     $n\sim\mathcal{N}(0,I)$
     $\tilde{g}(x)_{t}=I(x)(1-t)+\mathcal{Z}t+\epsilon\cdot n\cdot t$
     Take a gradient descent step on
$\nabla_{\theta}\|f_{\theta}(\tilde{g}(x)_{t},x,t)-I(x)\|_{2}$
until converged
@@ 3.2 Iterative Implicit Neural Representations: I-INRs
$\mathcal{Z}$, $f_{\theta}(\cdot,t)$, $\delta$, $x$
$\hat{g}(x)_{1}=\mathcal{Z}$
for $t\in\{1,1-\delta,1-2\delta,\dots,0\}$ do $\triangleright$ $t$ decreases from $1$ to 00 in steps of $\delta$
     $\hat{g}(x)_{t-\delta}=\frac{\delta}{t}f_{\theta}(\hat{g}(x)_{t},{x},t)+\left(1-\frac{\delta}{t}\right)\hat{g}(x)_{t}$
end for
return $\tilde{g}(x)_{0}$
Figure 4: Visual quality comparison of super-resolution results at 2× scale using various methods. The ground truth is compared against baseline INR methods SIREN [39], WIRE [36], Gauss [30], and their iterative counterparts. The iterative approaches consistently achieve sharper reconstructions with fewer artifacts, effectively preserving finer details and high-frequency structures.
Network Architecture: Figure 2(a) illustrates the I-INR architecture that integrates an Auxiliary Network and a Backbone (such as SIREN) to compute the output $I(x)_{t}$. The input $x$ is processed independently by both the Auxiliary Network and the Backbone, while the Auxiliary Network also incorporates feedback from the previous step. The output of the Auxiliary Network is concatenated with the features extracted by the Backbone, forming a joint representation. This concatenated feature then passes through a projection layer, which further refines the combined features. The final output is obtained as the dot product of the projection layer and the Backbone.

### v3

@@ Abstract
Implicit Neural Representations (INRs) have revolutionized signal processing and computer vision by modeling signals as continuous, differentiable functions parameterized by neural networks. However, INRs are prone to the spectral bias problem, limiting their ability to retain high-frequency information, and often struggle with noise robustness. Motivated by recent trends in iterative refinement processes, we propose Iterative Implicit Neural Representations (I-INRs). This novel plug-and-play framework iteratively refines signal reconstructions to restore high-frequency details, improve noise robustness, and enhance generalization, ultimately delivering superior reconstruction quality. I-INRs integrate seamlessly into existing INR architectures with only a 0.5–2% increase in parameters. During reconstruction, the iterative refinement adds just 0.8–1.6% additional FLOPs over the baseline while delivering a substantial performance boost of up to +2.0 PSNR. Extensive experiments demonstrate that I-INRs consistently outperform WIRE, SIREN, and Gauss across various computer vision tasks, including image fitting, image denoising, and object occupancy prediction. The code is available at https://github.com/optimizer077/I-INR.
[…]
@@ 1 Introduction
Implicit Neural Representations (INRs) have emerged as a transformative approach in signal representation, shifting from traditional discrete grid-based methods to continuous coordinate-based models (sitzmann2020implicit). By leveraging neural networks, typically multi-layer perceptrons (MLPs), INRs map spatial or temporal coordinates directly to signal attributes such as pixel intensity, colour, or 3D occupancy. This continuous representation inherently offers resolution independence, compact encoding, and seamless interpolation, making INRs highly versatile across a wide range of domains, including computer vision (chen2021learning; saragadam2022miner; sitzmann2020implicit; mildenhall2021nerf; jiang2023coordinate), and beyond (sun2021coil; shen2021non; zhong2019reconstructing; reed2021implicit). Initially introduced for tasks like 3D shape reconstruction (mildenhall2021nerf; peng2020convolutional; srinivasan2021nerv) and novel view synthesis (Irshad_2023_ICCV), INRs have since been applied to diverse applications including image fitting (liu2023finer), super-resolution (delbracioinversion; saharia2022image), and solving inverse problems (cha2024descanning; wang2019spatial; chen2021learning).
Despite advancements, INRs continue to face substantial challenges in capturing high-frequency details, maintaining robustness against noise, and handling incomplete data (saragadam2023wire). Typically optimized with L1 or L2 loss (sitzmann2020implicit), INRs inherently exhibit a spectral bias that favors low-frequency components, often at the expense of fine-grained details (rahaman2019spectral).
To mitigate this, recent approaches have introduced positional encoding schemes (tancik2020fourier; fathony2020multiplicative) and alternative activation functions (sitzmann2020implicit; saragadam2023wire; liu2023finer; gao2024h; thennakoon2025bandrc) to better capture high-frequency content. Positional encodings inject high-frequency signals through orthogonal Fourier bases (tancik2020fourier), while periodic activations such as sine functions enable MLPs to model high-frequency structures more effectively compared to traditional ReLU-based networks (sitzmann2020implicit). Nevertheless, these methods often assume clean and fully observed inputs, limiting their effectiveness in practical scenarios characterized by noise and occlusion (saragadam2023wire). More recent efforts, such as the integration of complex Gabor wavelet activations, have aimed to improve robustness under such conditions (saragadam2023wire), but challenges persist when scaling to real-world, noisy data. Consequently, there remains significant potential to enhance the fidelity, robustness, and generalization of INRs, particularly in preserving high-frequency signals under adverse conditions.
@@
To address these limitations, and drawing inspiration from iterative models (delbracioinversion; rissanen2022generative; ho2020denoising), we introduce Iterative Implicit Neural Representations (I-INRs) a plug-and-play framework that integrates seamlessly with existing implicit architectures. Unlike traditional single-shot INRs, I-INRs reconstruct signals progressively over multiple iterations, refining details, enhancing reconstruction quality, and improving noise robustness, as illustrated in Figure 1.
The proposed framework features a novel architecture that incorporates existing implicit networks as a BackboneNet without major modifications, while introducing two lightweight modules, i.e FeedbackNet and FuseNet as shown in Figure 2(a). These blocks add only 0.5–2% to the total parameter count of the baseline model and support the iterative reconstruction process. Owing to this design, the iterative I-INR process incurs only 0.8–2% additional FLOPs over the baseline INR for two-step reconstruction. Our experiments demonstrate that I-INRs deliver superior reconstructions with enhanced detail, better generalization, and minimal computational overhead, consistently outperforming single-shot methods such as WIRE (saragadam2023wire), SIREN (sitzmann2020implicit), and Gauss (ramasinghe2022beyond) across tasks including image fitting, super-resolution, denoising, and 3D occupancy reconstruction.
Our main contributions are as follows:
@@ 2 Related Work
Spatial Encoding Techniques.
INRs suffer from spectral bias (rahaman2019spectral; tancik2020fourier), struggling to capture high-fidelity details of complex signals. This limitation affects applications requiring fine-grained detail, such as a single image super resolution (SR), 3D reconstruction, and scene modeling. To address this challenge, various approaches have been proposed, including spatial encoding with frequency-based representations, polynomial decomposition (raghavan2023neural; singh2023polynomial), and high-pass filtering (fathony2020multiplicative), all of which emphasize high-frequency components. DINER (xie2023diner) applies a hash map to unevenly map input coordinates to feature vectors, optimizing spatial frequency distribution for faster, more accurate reconstructions.
Role of Activations.
Beyond encoding, activation functions critically influence spectral bias. Standard choices like ReLU cannot capture high-frequency components due to limited representational capacity (sitzmann2020implicit). To address this, alternatives such as sinusoidal (sitzmann2020implicit) and Gaussian (ramasinghe2022beyond) activations have been proposed. Activations tuned for high accuracy can suffer from reduced noise robustness; WIRE  (saragadam2023wire) leverages Gabor wavelets to provide enhanced robustness to noise. However, there remains room for improvement in achieving both high-frequency fitting and noise robustness  (saragadam2023wire).
Iterative Methods.
Iterative models have been very successful in image restoration and translation (saharia2022image; rissanen2022generative; delbracioinversion; hui2024microdiffusion; chu2025highly; chen2024image), video generation (ji2025layerflow), language modeling (nie2025large), 3D generation (qian2023magic123; liu2023zero), and more. In contrast with single-shot models, iterative models such as diffusion models (ho2020denoising; rissanen2022generative; delbracioinversion) produce high-quality samples by reversing degradation processes over multiple steps. Despite their widespread adoption and effectiveness, iterative models remain underexplored in the context of implicit neural representations.
@@
Inspired by the strong detail reconstruction capabilities of iterative models, we propose a novel INR framework that reconstructs signals implicitly over multiple steps. Our approach provides improved details and enhanced noise robustness, offering a flexible, plug-and-play solution that can be integrated into any existing INR framework. However, for the scope of this paper, we specifically focus on WIRE (saragadam2023wire), Gauss (ramasinghe2022beyond), and SIREN (sitzmann2020implicit).
@@
Figure 2: (a) The proposed architecture of the I-INR model. The framework consists of a Backbone, which can be any baseline INR architecture. Additionally, a Feedback Net incorporates feedback to refine representations, while FuseNet integrates features. The final output is obtained by combining the outputs of the Backbone and the FuseNet, enhancing expressivity and reconstruction quality. (b) Iterative reconstruction process of the proposed I-INR framework.
3 Preliminary
3.1 Standard INR
Let $g:X\subseteq\mathbb{R}^{p}\rightarrow Y\subseteq\mathbb{R}^{q}$ be a target function mapping $p$-dimensional inputs (such as pixel positions in a 2D image or volumetric domains in 3D scenes) to $q$-dimensional outputs (such as color or density values), such that $y=g(x)$ for $x\in X$. An INR (sitzmann2020implicit) is a learnable neural function $f_{\theta}:X\rightarrow Y$, parameterized by $\theta$, aiming to approximate $\mathcal{I}(x)$ across $X$. The objective is to find $\theta$ such that:
$f_{\theta}(x)\approx\mathcal{I}(x),$ (1)
where $\mathcal{I}(x)$ represents a true image function evaluated at coordinates $x$. A typical INR is trained by reducing the signal error in an $L_{p}$ metric:
$\min_{\theta}\mathbb{E}_{x}\left[\|f_{\theta}(x)-\mathcal{I}(x)\|_{p}\right]\approx\min_{\theta}\sum_{i}\|f_{\theta}(x^{i})-\mathcal{I}(x^{i})\|_{p},$ (2)
where $p=2$ for standard INR training, corresponding to the mean squared error.
3.2 Inversion by Direct Iteration (InDI)
Inversion by Direct Iteration (InDI) (delbracioinversion) addresses ill-posed inverse problems, such as image restoration, by iteratively refining an estimate of the unknown signal $b$ from an observed measurement $a$. Instead of solving the full inverse problem in one step, InDI interpolates between $a$ and $b$ via $b_{t}=ta+(1-t)b$, with $t\in[0,1]$, and uses a neural network $F_{\theta}$ to approximate the conditional expectation $\mathbb{E}[b\mid b_{t}]$. The network is trained to minimize:
$\min_{\theta}\,\mathbb{E}_{a,b,t,n}\left[\,\|F_{\theta}(b_{t}+\epsilon_{t}n,\,t)-b\|_{1}\,\right],$
where $\epsilon_{t}n$ is Gaussian noise. At inference, the estimate is updated recursively using the learned conditional mean, gradually solving a sequence of easier inverse problems from $a$ to $b$.
@@ 4 Proposed Method
In this section, we present a detailed explanation of our proposed methods, including a comprehensive description of the iterative plug-and-play framework.
4.1 Iterative Implicit Neural Representations: I-INR
I-INR reconstructs the signal over multiple steps, progressively improving the quality at each iteration. As shown in Figure 2(b), the process starts from an initial state $g(x)_{1}=\mathcal{Z}$ at $t=1$ and gradually approaches the final reconstruction $g(x)_{0}=\mathcal{I}(x)$ at $t=0$, using steps of size $\delta$. This is achieved by inverting the forward process:
$g(x)_{t}=\mathcal{I}(x)(1-t)+\mathcal{Z}t,\qquad t\in[0,1],$ (3)
where $\mathcal{Z}$ is sampled from a known distribution matching the shape of $\mathcal{I}(x)$.
At each step, I-INR estimates the next state by computing the conditional expectation $\mathbb{E}[g(x)_{0}\mid g(x)_{t}]$, which represents the best possible prediction of the original signal given the current state. This estimate is then used to update the reconstruction as follows:
$\hat{g}(x)_{t-\delta}$ $=\mathbb{E}\left[g(x)_{t-\delta}\mid\hat{g}(x)_{t}\right]$ (4)
$=\frac{\delta}{t}\,\mathbb{E}\left[g(x)_{0}\mid\hat{g}(x)_{t}\right]+\left(1-\frac{\delta}{t}\right)\hat{g}(x)_{t},$
by repeatedly applying this update, I-INR smoothly transitions from the initial state to the reconstructed signal, leveraging the conditional expectation at each step to improve accuracy.
I–INR Training.
To realize the update rule in Eq. (4), we train an implicit neural network $f_{\theta}$ (see Figure. 2(a)) that maps the intermediate state $\tilde{g}(x)_{t}$, the spatial coordinates $x$, and the time index $t$ directly to the clean target $\mathcal{I}(x)$:
$\min_{\theta}\;\mathbb{E}_{x,\,t,\,n}\left\|f_{\theta}\!\bigl(\tilde{g}(x)_{t},\,x,\,t\bigr)-\mathcal{I}(x)\right\|_{2}^{2},$ (5)
$\tilde{g}(x)_{t}=(1-t)\,\mathcal{I}(x)+t\,\mathcal{Z}+\varepsilon tn,$ (6)
where $\mathcal{Z}\sim p(\mathcal{Z})$ and $n\sim\mathcal{N}(0,1)$. The small perturbation $\varepsilon tn$ (delbracioinversion) satisfies the regularity requirement, guaranteeing stable reconstruction during inference. The full training procedure appears in Algorithm 1.
I–INR Reconstruction.
Reconstruction begins from the latent state $\mathcal{Z}$ and proceeds for $1/\delta$ iterations of Eq.  4, iteratively refining the estimate towards the final signal (Algorithm 2).
Algorithm 1 I-INRs Model Training of $f_{\theta}$
@@ 4.1 Iterative Implicit Neural Representations: I-INR
[…]
3:repeat
4:  $x\sim p(x)$
5:  $t\sim\mathcal{U}(0,1),\;n\sim\mathcal{N}(0,I)$
6:  $\tilde{g}(x)_{t}\leftarrow(1-t)\,\mathcal{I}(x)+t\,\mathcal{Z}+\varepsilon\,t\,n$
7:  $\theta\leftarrow\theta-\eta\,\nabla_{\theta}\bigl\|f_{\theta}(\tilde{g}(x)_{t},\,x,\,t)-\mathcal{I}(x)\bigr\|_{2}^{2}$
8:until converged
@@
[…]
3:for $t\leftarrow 1$ to 0 step $\delta$ do
4:  $\hat{g}(x)_{t-\delta}\leftarrow\dfrac{\delta}{t}\,f_{\theta}\bigl(\hat{g}(x)_{t},x,t\bigr)+\left(1-\dfrac{\delta}{t}\right)\hat{g}(x)_{t}$
5:end for
6:return $\hat{g}(x)_{0}$
[…]
Network Architecture.
Figure 2(a) provides an overview of our plug-and-play I-INR framework, which comprises three main components: a Backbone network, a FeedbackNet module, and a FuseNet module. The Backbone extracts initial features from the input; the FeedbackNet incorporates the intermediate state and time conditioning; and the FuseNet merges these feature streams. The final output is given by Eq. (7). This design enables flexible integration with existing INR architectures and supports efficient iterative refinement. The information flow between components is formalized as
$f_{\theta}\bigl(\hat{g}(x)_{t},\,x,\,t\bigr)=\operatorname{FuseNet}\bigl(\operatorname{concat}(\mathbf{f},\mathbf{b})\bigr)\odot\mathbf{b},$ (7)
where
$\mathbf{b}:=\operatorname{Backbone}(x),\qquad\mathbf{f}:=\operatorname{FeedbackNet}\bigl(\hat{g}(x)_{t},\,x,\,t\bigr).$
At each iteration, the current state is processed by the FeedbackNet, and its output is fused with the Backbone features via FuseNet. Notably, the Backbone is forward-passed only once throughout the reconstruction, while the lightweight FeedbackNet and FuseNet operate at each iteration, resulting in a substantial reduction in computational cost.